OpenStudy (anonymous):

write each quotient in simplest terms. 3/11 divided by 6/7

OpenStudy (anonymous):

\[3/11*7/6=(7*3)/(11*6)\]

OpenStudy (anonymous):

3/11*7/6 = 21/66 =7/22 ooh inverse PI

OpenStudy (anonymous):

They are able to be reduced by a factor of 3